# High dimensional hyperbolic Coxeter groups that virtually fiber

### v1

@@ 1 Introduction
We now describe the construction. In [12], Osajda developed a recursive construction that builds hyperbolic right-angled Coxeter groups with arbitrarily large virtual cohomological dimension. In general, these Coxeter groups may not satisfy the hypotheses of [10, Thm. 6.14], and therefore the combinatorial systems of moves approach developed in that paper may not apply to provide a virtual algebraic fibration. We combine the recursive construction in [12] with a novel “$\alpha$-thickening” procedure generalizing the thickening procedure used by Osajda to construct hyperbolic right-angled Coxeter groups whose defining graphs admit a state with a legal orbit under an appropriate system of moves, so [10] then applies to provide a virtual fibration.
Given a finite cube complex $X$, our $\alpha$-thickening procedure enlarges the vertex set by a surjection $\alpha:Y\to V(X)$ from a large finite set $Y$, inserts a high-dimensional simplex above each vertex of the original cube complex, and then applies the standard thickening procedure of [12]. This has the effect of inserting enough vertices and edges into the 1-skeleton of the complex so that one can develop a legal system of moves without changing the homotopy type of the resulting simplicial complex. This process may be of independent interest and is developed in Section 2.2 below.
@@ Acknowledgements
This work was completed as part of the ‘Collaborative Projects in Geometric Topology’ program organized by the Geometry and Topology group at the University of Virginia and supported by an RTG grant from the National Science Foundation (DMS-1839968). We thank the main organizers Sara Maloni and Thomas Mark for providing us the opportunity to work together on this project. We also thank Mike Davis, Jingyin Huang, and Boris Okun for some helpful comments and references. Lafont and Stover were partially supported by the NSF under grants DMS-2407438 and DMS-2203555, respectively.
@@ 2.1 Jankiewicz–Norin–Wise States and Moves
Let $\Gamma=\Gamma(V)$ be a finite simplicial graph with vertex set $V$. A state of $\Gamma$ is a function ${\varphi:V\to\mathbb{Z}_{2}}$, which amounts to an assignment of 0 or 1 to every vertex in the graph. A state is said to be legal if the subgraphs of $V$ induced by $\varphi^{-1}(0)$ and $\varphi^{-1}(1)$ are both connected and nonempty. A move at $v\in V$ is a state $m_{v}$ with the following properties:
@@ 2.2 Simplicial Thickenings of Cube Complexes
We recall some terminology related to simplicial complexes. A subcomplex $\Sigma\prime\subset\Sigma$ is full if every simplex of $\Sigma$ whose vertices lie in $\Sigma\prime$ is actually contained in $\Sigma\prime$. A simplicial complex $\Sigma$ is flag if every $(k+1)$-tuple of pairwise adjacent vertices spans a $k$-simplex. Flag simplicial complexes are completely determined by their $1$-skeleton. For $k\geq 5$, we say a simplicial complex $\Sigma$ is $k$-large if it is flag, and any embedded cycle of length $<k$ fails to be full. Note that, for $k=5$, this implies that for any embedded $4$-cycle $(v_{1},v_{2},v_{3},v_{4})$, one of the diagonals must be present, i.e. either $v_{1}$ is joined to $v_{3}$ by an edge, or $v_{2}$ is joined to $v_{4}$. The property of being $5$-large is sometimes also called the “flag-no-square” condition.
@@
Two distinct vertices $v$ and $w$ in a cube complex $X$ have cubical distance 1 if and only if they are contained in a common cube. More generally, the cubical distance $d_{\text{cube}}(v,w)$ between distinct vertices $v$ and $w$ is the minimum length of a sequence of vertices ${v=v_{1},\ldots,v_{n}=w}$ where $d_{\text{cube}}(v_{i},v_{i+1})=1$ for each $i$.
@@
Let $X$ be a finite cube complex with vertex set $V(X)$. Given a finite set $Y$ equipped with a surjective map $\alpha:Y\to V(X)$, the $\alpha$-thickening of $X$ is the unique simplicial complex $\operatorname{Th}_{\alpha}(X)$ with vertex set $Y$ so that $y_{0},\dots,y_{k}\in Y$ span a $k$-simplex of $\operatorname{Th}_{\alpha}(X)$ if and only if $\alpha(y_{0}),\dots,\alpha(y_{k})$ are contained in a common cube of $X$.
@@
  • If $X=\{x\}$ and $Y=\{y_{0},\ldots,y_{n}\}$, then $\alpha:Y\to X$ is the constant map and $\operatorname{Th}_{\alpha}(X)$ is a single $n$-simplex.
  • On any cube complex $X$, we can recover the thickening $\operatorname{Th}_{1}(X)$ defined by Osajda [12], as it coincides with the $\alpha$-thickening associated with the identity map ${\alpha=\mathrm{Id}:V(X)\to V(X)}$.
We also note that there is an embedding of the Osajda thickening $\operatorname{Th}_{1}(X)\hookrightarrow\operatorname{Th}_{\alpha}(X)$ into the $\alpha$-thickening associated with any section of the map $\alpha:Y\rightarrow V(X)$. Moreover, $\operatorname{Th}_{\alpha}(X)$ deformation retracts to the image of $\operatorname{Th}_{1}(X)$, and hence we have a homotopy equivalence $\operatorname{Th}_{\alpha}(X)\simeq\operatorname{Th}_{1}(X)$. Combining this with the homotopy equivalence $\operatorname{Th}_{1}(X)\simeq X$ (see [12, Lem. 3.5]) we see that $\operatorname{Th}_{\alpha}(X)\simeq X$. Next we show that the thickening inherits $5$-largeness.
@@
Suppose $X$ is a $5$-large cubical complex. Then for any surjective $\alpha:Y\to V(X)$, the simplicial complex $\operatorname{Th}_{\alpha}(X)$ is 5-large.
@@
Consider the universal cover $\tilde{X}$ of $X$. Since $\tilde{X}$ is a locally $5$-large, simply connected cube complex, we have that $\operatorname{Th}_{1}(\tilde{X})$ is $5$-large; see [12, Prop. 3.4]. The covering map $\tilde{X}\rightarrow X$ naturally induces a covering map $\operatorname{Th}_{1}(\tilde{X})\rightarrow\operatorname{Th}_{1}(X)$. Our hypothesis guarantees that cubical $2$-neighborhoods in $X$ lift to cubical $2$-neighborhoods in $\tilde{X}$. It follows that the combinatorial $2$-neighborhoods in $\operatorname{Th}_{1}(X)$ lift isomorphically to combinatorial $2$-neighborhoods in $\operatorname{Th}_{1}(\tilde{X})$. However, the properties of being flag and $5$-large are local: they only involve cycles of length at most $4$. Any such cycle is contained in a combinatorial $2$-neighborhood, so $\operatorname{Th}_{1}(X)$ is also $5$-large. For the general case where $\operatorname{Th}_{\alpha}(X)$ is an arbitrary thickening, it is straightforward to use the embeddings $\operatorname{Th}_{1}(X)\hookrightarrow\operatorname{Th}_{\alpha}(X)$ to reduce to the $\operatorname{Th}_{1}(X)$ case. We leave the details to the reader. ∎
@@ 3.1 Thickening $\mathbb{T}$, states, and moves
We begin by defining the thickening $\mathbb{T}(X)$ of a cube complex $X$ (shown in Figure 2) that, when applied to a cube complex $X$ with the appropriate properties, will admit a legal system of moves.
@@
The vertices of $\mathbb{T}(X)$ are defined to be the following subset of $V(X)\times V(X)$:
$V(\mathbb{T}(X))=\left\{(v,w)\;:\;d_{\text{cube}}(v,w)\geq 2\,\text{ in }X\right\}$
The map $\alpha:V(\mathbb{T}(X))\rightarrow V(X)$ is simply projection onto the first factor.
@@
Suppose $v,w\in V(X)$ with $d_{\text{cube}}(v,w)\geq 2$, so there are associated vertices $v_{w},w_{v}\in\mathbb{T}(X)$. The key property of our starting state $\mathcal{S}$ is that it assigns opposite values to these two vertices. If $v=x_{i}$, $w=x_{j}$, and $i<j$, then $\mathcal{S}(v_{w})=1$ and $\mathcal{S}(w_{v})=0$.
We now define the system of moves that, when paired with the state $\mathcal{S}$ from equation (1), produce a legal system on $\mathbb{T}(X)$.
@@
Given any pair of vertices $v,w\in X$ at cubical distance $\geq 2$, we define the function $m_{vw}:V(\mathbb{T}(X))\rightarrow\mathbb{Z}_{2}$ by setting $m_{vw}(v_{w})=m_{vw}(w_{v})=1$ and $m_{vw}$ to be identically zero on all other vertices.
Note that the vertices $v_{w}$ and $w_{v}$ are not adjacent in $\mathbb{T}(X)$, since they are at cubical distance at least $2$ in $X$. This implies that each function $m_{vw}$ is a move at the vertices $v_{w}$ and $w_{v}$ (and in fact, $m_{vw}=m_{wv}$). Informally, the move $m_{vw}$ toggles the $\mathbb{Z}_{2}$-assignment of both vertices $v$ and $w$, and preserves the $\mathbb{Z}_{2}$-assignment of all other vertices. Let $M$ denote the subgroup generated by the system of moves.
@@ 3.2 Verifying the system is legal
The thickening $\mathbb{T}(X)$ can be applied to any finite cube complex $X$. We will now find suitable conditions on $X$ to guarantee the initial state and moves on $\mathbb{T}(X)$ provide a legal system. The required assumptions on $X$ are contained in the next two definitions.
@@
A cube complex $X$ has no isolated corners if, given any cube $\square\subset X$ and vertex $v\in\square$, there exists a vertex $w$ adjacent to $v$ and satisfying $w\notin\square$. It has no disconnecting cubes if there is no cube whose removal disconnects $X$.
Our goal in this section will be to establish the following
@@
$\delta_{\epsilon}:V(X)\longrightarrow\{\text{Y},\text{N}\}$
@@
The assumptions that there are no disconnecting cubes and no isolated corners on $X$ are not only sufficient for $X$ to admit an $\alpha$-thickening with a legal system of moves, but they are also necessary. Suppose $X$ does not satisfy the no disconnecting cube condition, and let $\square$ denote a maximal cube whose removal disconnects $X$. Then one can apply a sequence of moves making every vertex $v\in\square$ satisfy $\delta_{\epsilon}(v)=\text{N}$. In this way, $\square$ would serve as an “$\epsilon$-blockade” between the two components of $X\setminus\square$, and thus the corresponding system of moves would not be legal. If $X$ does not satisfy the no isolated corner condition, then any such corner vertex $v\in\square$ could be similarly blocked from the rest of the cube complex, by applying a sequence of moves that change the vertices above all the vertices in $\square\setminus v$.
@@ 4 Proof of the Main Theorem
$\operatorname{vcd}(G_{n+1})=\max_{\sigma\subset T_{n}}\!\left\{n\ :\ \overline{H}^{n-1}(T_{n}\setminus\sigma)\neq 0\right\}\!.$ (2)
@@
Here the cohomological dimension $\operatorname{cd}(Z)$ of a space $Z$ is the largest integer $k$ such that $H^{k}(Z)$ is nontrivial, which is a slight abuse of terminology [^1] This is not the standard definition, as cohomological dimension is usually defined as the largest $k$ so that $H^{k}(Z,C)\neq 0$ for some closed subset $C\subset Z$. Throughout our paper we are working with the special case where $C=\emptyset$. Nevertheless, it will be convenient to use the notation $\operatorname{cd}(Z)$, which could be smaller than the classically defined cohomological dimension of $Z$. We hope this will not cause confusion for the reader.. The proof of 4.1 will be postponed until the next section, as it is somewhat involved. Assuming this result, we can now continue our verification of the second step.
@@
The group $G_{n+1}$ has virtual cohomological dimension $=n+1$.
@@
To show that $\operatorname{vcd}(G_{n+1})=n+1$ we must verify that $\operatorname{cd}(T_{n}\setminus\sigma)\leq n$ for each nonempty simplex $\sigma\subset T_{n}$. Given such a simplex $\sigma$, cover $T_{n}$ by two open sets $U,V$, by setting $U$ to be an $\epsilon$-neighborhood of $\sigma$, and $V$ to be the complement $T_{n}\setminus\sigma$. The intersection $U\cap V$ is then homotopy equivalent to $S(\sigma,T_{n})$. Since $U$ is contractible, the Mayer–Vietoris sequence in cohomology reduces to the long exact sequence:
$\cdots\longrightarrow H^{i}(T_{n})\longrightarrow H^{i}(T_{n}\setminus\sigma)\longrightarrow H^{i}(S(\sigma,T_{n}))\longrightarrow H^{i+1}(T_{n})\longrightarrow\cdots$
Since $\operatorname{cd}(T_{n})=n$, the sequence gives us isomorphisms $H^{i}(T_{n}\setminus\sigma)\cong H^{i}(S(\sigma,T_{n}))$ in all degrees $i\geq n+1$. However we know from 4.1 that $\operatorname{cd}(S(\sigma,T_{n}))\leq n-1$, so the groups $H^{i}(T_{n}\setminus\sigma)$ then vanish when $i\geq n+1$. We conclude that $\operatorname{cd}(T_{n}\setminus\sigma)\leq n$, as desired. This completes the proof of the proposition. ∎
@@
For a suitable finite index torsion-free subgroup $\Lambda\leq G_{n+1}$, the finite cube complex $X_{n+1}\coloneqq D_{n+1}/\Lambda$ is $5$-large, has no disconnecting cubes, has no isolated corners, and has cohomological dimension $=n+1$.

### v2

@@ 1 Introduction
We now describe the construction. In [12], Osajda developed a recursive construction that builds hyperbolic right-angled Coxeter groups with arbitrarily large virtual cohomological dimension. In general, these Coxeter groups may not satisfy the hypotheses of [10, Thm. 6.14], and therefore the combinatorial systems of moves approach developed in that paper may not apply to provide a virtual algebraic fibration. We combine the recursive construction in [12] with a novel $\alpha$-thickening” procedure generalizing the thickening procedure used by Osajda to construct hyperbolic right-angled Coxeter groups whose defining graphs admit a state with a legal orbit under an appropriate system of moves, so [10] then applies to provide a virtual fibration.
Given a finite cube complex $X$, our $\alpha$-thickening procedure enlarges the vertex set by a surjection $\alpha\colon Y\to V(X)$ from a large finite set $Y$, inserts a high-dimensional simplex above each vertex of the original cube complex, and then applies the standard thickening procedure of [12]. This has the effect of inserting enough vertices and edges into the 1-skeleton of the complex so that one can develop a legal system of moves without changing the homotopy type of the resulting simplicial complex. This process may be of independent interest and is developed in Section 2.2 below.
@@ Acknowledgements
This work was completed as part of the ‘Collaborative Projects in Geometric Topology’ program organized by the Geometry and Topology group at the University of Virginia and supported by an RTG grant from the National Science Foundation (DMS-1839968). We thank the main organizers Sara Maloni and Thomas Mark for providing us the opportunity to work together on this project. We also thank Mike Davis, Jingyin Huang, and Boris Okun for some helpful comments and references. Lafont was partially supported by the NSF under grant DMS-2407438. Stover was partially supported by NSF grants DMS-2203555 and DMS-2506896, along with award SFI-MPS-TSM-00014184 from the Simons Foundation. The authors thank the referee for a very careful reading of the paper.
@@ 2.1 Jankiewicz–Norin–Wise States and Moves
Let $\Gamma=\Gamma(V)$ be a finite simplicial graph with vertex set $V$. A state of $\Gamma$ is a function $\varphi\colon V\to\mathbb{Z}_{2}$}, which amounts to an assignment of 0 or 1 to every vertex in the graph. A state is said to be legal if the subgraphs of $V$ induced by $\varphi^{-1}(0)$ and $\varphi^{-1}(1)$ are both connected and nonempty. A move at $v\in V$ is a state $m_{v}$ with the following properties:
@@ 2.2 Simplicial Thickenings of Cube Complexes
We recall some terminology related to simplicial complexes. A subcomplex $\Sigma\prime\subset\Sigma$ is full if every simplex of $\Sigma$ whose vertices lie in $\Sigma\prime$ is actually contained in $\Sigma\prime$. A simplicial complex $\Sigma$ is flag if every $(k+1)$-tuple of pairwise adjacent vertices spans a $k$-simplex. Flag simplicial complexes are completely determined by their $1$-skeleton. For $k\geq 5$, we say a simplicial complex $\Sigma$ is $k$-large if it is flag, and any embedded cycle of length <k<k fails to be full. Note that, for $k=5$, this implies that for any embedded $4$-cycle $(v_{1},v_{2},v_{3},v_{4})$, one of the diagonals must be present, i.e. either $v_{1}$ is joined to $v_{3}$ by an edge, or $v_{2}$ is joined to $v_{4}$. The property of being $5$-large is sometimes also called the “flag-no-square” condition.
@@
Two distinct vertices $v$ and $w$ in a cube complex $X$ have cubical distance 1 if and only if they are contained in a common cube. More generally, the cubical distance $d_{\text{cube}}(v,w)$ between distinct vertices $v$ and $w$ is the minimum number $n$ such that there is a sequence of vertices $v=v_{0},\ldots,v_{n}=w$} where $d_{\text{cube}}(v_{i},v_{i+1})=1$ for each $i$.
@@
Let $X$ be a finite cube complex with vertex set $V(X)$. Given a finite set $Y$ equipped with a surjective map $\alpha\colon Y\to V(X)$, the $\alpha$-thickening of $X$ is the unique simplicial complex $\operatorname{Th}_{\alpha}(X)$ with vertex set $Y$ so that $y_{0},\dots,y_{k}\in Y$ span a $k$-simplex of $\operatorname{Th}_{\alpha}(X)$ if and only if the images $\alpha(y_{0}),\dots,\alpha(y_{k})$} are contained in a common cube of $X$.
@@
  • If $X=\{x\}$ and $Y=\{y_{0},\ldots,y_{n}\}$, then $\alpha\colon Y\to X$ is the constant map and $\operatorname{Th}_{\alpha}(X)$ is a single $n$-simplex.
  • On any cube complex $X$, we can recover the thickening $\operatorname{Th}_{1}(X)$ defined by Osajda [12], as it coincides with the $\alpha$-thickening associated with the identity map $\alpha=\mathrm{Id}\colon V(X)\to V(X)$}.
We also note that there is an embedding of the Osajda thickening $\operatorname{Th}_{1}(X)\hookrightarrow\operatorname{Th}_{\alpha}(X)$ into the $\alpha$-thickening associated with any section of the map $\alpha\colon Y\rightarrow V(X)$. Moreover, $\operatorname{Th}_{\alpha}(X)$ deformation retracts to the image of $\operatorname{Th}_{1}(X)$, and hence we have a homotopy equivalence $\operatorname{Th}_{\alpha}(X)\simeq\operatorname{Th}_{1}(X)$. Combining this with the homotopy equivalence $\operatorname{Th}_{1}(X)\simeq X$ (see [12, Lem. 3.5]) we see that $\operatorname{Th}_{\alpha}(X)\simeq X$. Next we show that the thickening inherits $5$-largeness.
@@
Suppose $X$ is a $5$-large cubical complex. Then for any surjective $\alpha\colon Y\to V(X)$, the simplicial complex $\operatorname{Th}_{\alpha}(X)$ is 5-large.
@@
[…]
@@ 3.1 Thickening $\mathbb{T}$, states, and moves
We begin by defining a particular thickening $\mathbb{T}(X)$ of a cube complex $X$ (shown in Figure 2) that, when applied to a cube complex $X$ with the appropriate properties, will admit a legal system of moves.
@@
Let $X$ be a cube complex for which every vertex is at cubical distance $2$ from some other vertex. The thickening $\mathbb{T}(X)$ has vertices given by the following subset of $V(X)\times V(X)$:
[…]
The map $\alpha\colon V(\mathbb{T}(X))\rightarrow V(X)$ is simply the projection onto the first factor.
@@
Suppose $v,w\in V(X)$ with $d_{\text{cube}}(v,w)\geq 2$, so there are associated vertices $v_{w},w_{v}\in\mathbb{T}(X)$. The key property of our starting state $\mathcal{S}$ is that it assigns opposite values to these two vertices. If $v=x_{i}$, $w=x_{j}$, and $i<j$, then $\mathcal{S}(v_{w})=1$ and $\mathcal{S}(w_{v})=0$. We now define the system of moves that, when paired with the state $\mathcal{S}$ described in (1), produce a legal system on $\mathbb{T}(X)$.
@@
Given any pair of vertices $v,w\in X$ at cubical distance $\geq 2$, we define the function $m_{vw}\colon V(\mathbb{T}(X))\rightarrow\mathbb{Z}_{2}$ by setting $m_{vw}(v_{w})=m_{vw}(w_{v})=1$ and $m_{vw}$ to be identically zero on all other vertices.
Note that the vertices $v_{w}$ and $w_{v}$ are not adjacent in $\mathbb{T}(X)$, since they are at cubical distance at least $2$ in $X$. This implies that each function $m_{vw}$ is a move at the vertices $v_{w}$ and $w_{v}$ (and in fact, $m_{vw}=m_{wv}$). Informally, the move $m_{vw}$ toggles the $\mathbb{Z}_{2}$-assignment of both vertices $v_{w}$ and $w_{v}$, and preserves the $\mathbb{Z}_{2}$-assignment of all other vertices. Let $M$ denote the subgroup generated by the system of moves.
@@ 3.2 Verifying the system is legal
The thickening $\mathbb{T}(X)$ can be applied to any finite cube complex $X$ for which every vertex is at cubical distance $2$ from some other vertex. We will now find suitable conditions on $X$ to guarantee the initial state and moves on $\mathbb{T}(X)$ provide a legal system. The required assumptions on $X$ are contained in the next two definitions.
@@
A cube complex $X$ has no isolated corners if, given any cube $\square\subset X$ and vertex $v\in\square$, there exists a vertex $w$ adjacent to $v$ and satisfying $w\notin\square$. $X$ has no disconnecting cubes if, for every $\square\subset X$, the full subcomplex with vertex set $V(X)\setminus V(\square)$ is connected.
Our goal in this section will be to establish the following.
@@
$\delta_{\epsilon}\colon V(X)\longrightarrow\{\text{Y},\text{N}\}$
@@
As shown above, the assumptions that there are no disconnecting cubes and no isolated corners on $X$ are sufficient for $X$ to admit an $\alpha$-thickening with a legal system of moves. For large cube complexes, one would expect that these conditions are also necessary by the following reasoning. Suppose $X$ does not satisfy the no disconnecting cube condition, and let $\square$ denote a maximal cube whose removal disconnects $X$. Then one can apply a sequence of moves making every vertex $v\in\square$ satisfy $\delta_{\epsilon}(v)=\text{N}$. In this way, $\square$ would serve as an $\epsilon$-blockade” between the two components of $X\setminus\square$, and thus the corresponding system of moves would not be legal. If $X$ does not satisfy the no isolated corner condition, then any such corner vertex $v\in\square$ could be similarly blocked from the rest of the cube complex, by applying a sequence of moves that change the vertices above all the vertices in $\square\setminus v$. For small cube complexes (say, for example, three vertices connected as a line segment), these assumptions are not quite necessary, but one would expect these conditions to be necessary for larger complexes.
@@ 4 Proof of the Main Theorem
$\operatorname{vcd}(G_{n+1})=\max\!\left\{k\ :\ \overline{H}{}^{k-1}(T_{n}\setminus\sigma)\neq 0\textrm{ for some }\sigma\subset T_{n}\right\}\!.$ (2)
@@
[…]
@@
The group $G_{n+1}$ has virtual cohomological dimension =n+1=n+1.
@@
To show that $\operatorname{vcd}(G_{n+1})=n+1$ we must verify that $\operatorname{cd}(T_{n}\setminus\sigma)\leq n$ for each nonempty simplex $\sigma\subset T_{n}$. Given such a simplex $\sigma$, cover $T_{n}$ by two open sets $U,V$, by setting $U$ to be an $\epsilon$-neighborhood of $\sigma$, and $V$ to be the complement $T_{n}\setminus\sigma$. The intersection $U\cap V$ is then homotopy equivalent to $S(\sigma,T_{n})$. Since $U$ is contractible, the Mayer–Vietoris sequence in reduced cohomology reduces to the long exact sequence:
[…]
@@
For a suitable finite index torsion-free subgroup $\Lambda\leq G_{n+1}$, the finite cube complex $X_{n+1}\coloneqq D_{n+1}/\Lambda$ is $5$-large, has no disconnecting cubes, has no isolated corners, and has cohomological dimension =n+1=n+1.
@@
Remark 4.5.
Note that the cube complex $X_{n+1}$ has the property that there exists another vertex at cubical distance $\geq 2$ from $v$ for every vertex $v$. This follows from the fact that $X_{n+1}$ is a quotient of the unbounded cube complex $D_{n+1}$ and that cubical $2$-neighborhoods in $X_{n+1}$ lift to cubical $2$-neighborhoods in $D_{n+1}$. This confirms that we can apply our thickening process to $X_{n+1}$ (see 3.1).